OpenStudy (anonymous):

Which of the following represents the linear equation 2x = y + 3 in slope−intercept form?

OpenStudy (anonymous):

u

OpenStudy (anonymous):

2x + y = 3 2x − y = 3 y = 2x − 3 y = 2x + 3